=================================================================== Tohban report: 2009/04/14 - 2009/04/21 last data checked 2009/04/19 00-24 UT =================================================================== Data Collection Status ======================= THEMIS A FGM: Full coverage until Apr 18 23 UT ESA: Full coverage until Apr 19 02 UT SST: Full coverage until Apr 19 02 UT EFI: Full coverage until Apr 19 01 UT SCM: Full coverage until Apr 15 01 UT THEMIS B FGM: data gap Apr 17 00-24 UT ESA: expected data gap from Apr 12 1600 UT - Apr 15 0430 UT (slow survey out coming out of perigee) data gap from Apr 15 1000 UT - Apr 16 1800 UT data gap from Apr 16 2300 UT - Apr 19 0700 UT SST: expected data gap from Apr 12 1600 UT - Apr 15 0430 UT (slow survey out coming out of perigee) data gap from Apr 15 1000 UT - Apr 16 1800 UT data gap from Apr 16 2300 UT - EFI: data gap Apr 17 00-24 UT SCM: data gap Apr 17 00-24 UT THEMIS c FGM: full coverage until Apr 18 0500 UT ESA: full coverage until Apr 18 0900 UT SST: full coverage until Apr 18 0900 UT EFI: full coverage until Apr 18 0900 UT SCM: full coverage until Apr 18 0900 UT THEMIS D FGM: full coverage until Apr 19 1200 UT ESA: full coverage until Apr 18 1600 UT data gap Apr 18 1600-2400 UT (data not dumped/processed yet) full coverage Apr 19 1600 UT SST: full coverage until Apr 19 1600 UT full coverage until Apr 15-11 1600 UT EFI: full coverage until Apr 19 1600 UT SCM: full coverage until Apr 19 1600 UT THEMIS E FGM: full coverage until Apr 19 1300 UT ESA: full coverage until Apr 19 1600 UT SST: full coverage until Apr 19 1600 UT EFI: full coverage until Apr 19 1600 UT SCM: full coverage until Apr 19 1600 UT Triggers and burst data: ========================= Inner probes (D, E, A) PB trigger Bz WB trigger filter banks FB1 = EFI, E12DC, 512-1024 Hz FB2 = SCM, SCM1, 64-128 Hz Outer probes (B,C in Dusk config): PB trigger on density WB trigger on filter banks FB1 = EFI, E12DC, 512-1024 Hz FB2 = SCM, SCM1, 1024-2048 Hz +ICCR 085 implemented on 4/13th on B and C (PB density trigger for dusk side magnetpause) Result:-Multiple burst captures on C on the 15th on sharp denisty gradients 0030 UT and 12 UT - trigger appears to be working. +ICCR 086 implemented on 4/16 23:41 on A. This has a new WB trigger for A FB1: EFI, input=E12DC, band=2048–4096 Hz, byte setting = 0x01 FB2: SCM, input=SCM1, band=16–32 Hz, byte setting = 0x06 to captured Kinetic Alfven waves and whistlers across magnetopause Result:-Wave burst captures largely where expected, though triggering seems more sensitive to changes in the EFI field FBK than SCM FBK (see data from the 17th for example) +ICCRF_JSM087 implement 17:13 UT Apr 17 on D which sets IIR filtering for the Filter Bank to remove spikeness causing spurious triggers. Result: This change appear to working well with WB captures where expected. Comments --------- Electric field data on A is improving significantly - spikes are of much lower amplitude (sphere shadowing) Significant events ------------------- +Apr 14 BGS_11M THEMIS_E -9.200 2009/104 18:20:00 2009/104 18:53:00 TR10 # ---> Apogee raise maneuver [DP] - P4 SM1C_D2 - MAN ID 267 - nominal burn - # spin rate adjust maneuver [DP] - P4 - MAN ID 268 - target spin rate 19.7938+/- 0.0130 rpm - nominal burn MILA_9M2 THEMIS_E 9.200 2009/104 19:12:00 2009/104 19:32:00 TR9 # Data dump - ATS setup for TR2 - blind acquisition MILA_9M2 THEMIS_E 2.690 2009/104 20:00:00 2009/104 20:20:00 TR2 # SOH check and track [DP] - post-maneuver instrument operations - blind acquisition +Apr 15 BGS_11M THEMIS_D 9.300 2009/105 17:04:50 2009/105 17:31:10 TR10 # ---> Apogee raise maneuver [ME] - P3 SM1C_D2 - MAN ID 269 - nominal burn AGO_9M THEMIS_D 9.300 2009/105 17:47:00 2009/105 18:12:00 TR10 # ---> Spin rate adjust maneuver [ME] - P3 - MAN ID 270 - target spin rate 19.7938 +/- 0.0130 rpm - nominal burn BGS_11M THEMIS_E 9.200 2009/105 18:28:00 2009/105 18:46:30 TR10 # Data dump AGO_9M THEMIS_D 2.680 2009/105 19:20:00 2009/105 19:40:00 TR2 # SOH check and track [ME] - post-maneuver instrument operations +Apr 16 AGO_9M THEMIS_D 2.780 2009/106 19:20:00 2009/106 19:35:00 TR2 # SOH check and track [JM] - no real-time network connections due to T1 line outage at JPL - no data loss BGS_11M THEMIS_B 2.950 2009/106 21:12:05 2009/106 21:32:05 TR2 # SOH check and track [JM] - enabled compression of APID 0x440 packets - ICCR_xxx BGS_11M THEMIS_C 2.850 2009/106 22:28:00 2009/106 22:48:00 TR2 # SOH check and track [JM] - enabled compression of APID 0x440 packets - ICCR_xxx BGS_11M THEMIS_A 9.100 2009/106 23:41:30 2009/107 00:02:30 TR10 # Data dump [DP] - configured filter bank settings for dusk season - ICCR_DOP86 +Apr 17 BGS_11M THEMIS_D 9.300 2009/107 16:57:30 2009/107 17:22:30 TR10 # Data dump [JM] - enable IIR filtering on the filter bank - ICCR_JSM87 BGS_11M THEMIS_E 9.200 2009/107 18:01:00 2009/107 18:36:15 TR10 # Data dump - blind acquisition - ATS turns transmitter on at 18:20:40 BGS_11M THEMIS_A 9.100 2009/108 00:12:30 2009/108 00:32:10 TR10 # Data dump [GL] - pre-maneuver table loads - instrument safing via A +Apr 18 TDRS_3 THEMIS_A 9.300 2009/109 00:05:00 2009/109 00:21:00 TR2 # Proficiency pass [ME] - planned maneuver not executed - hand-over to TDRS 6 - TDRS_6 THEMIS_A 9.300 2009/109 00:21:30 2009/109 00:40:00 TR2 # hand-over from TDRS 3 - hand-over to BGS BGS_11M THEMIS_A -9.100 2009/109 00:42:00 2009/109 02:01:00 TR9 # Data dump [ME] - post-maneuver instrument operations +Apr 19 BGS_11M THEMIS_B 9.500 2009/110 08:41:00 2009/110 09:26:00 TR19 # Data dump - pass starts during shadow BGS_11M THEMIS_C 9.500 2009/110 09:31:00 2009/110 10:16:00 TR39 # Data dump - pass during post-shadow period Science ========= +Apr 13-18 A, D, E: multiple dipolarisations with burst captures +Apr 14 A: Plasma sheet oscillations 1000 UT +Apr 15 B: Magnetopause crossing with good burst capture